# Nonlinear Online Optimization for Vehicle-Home-Grid Integration including Household Load Prediction and Battery Degradation

### v2

@@ Abstract
This paper investigates the economic impact of vehicle-home-grid integration, by proposing an online energy management algorithm that optimizes energy flows between an electric vehicle (EV), a household, and the electrical grid. The algorithm leverages vehicle-to-home (V2H) for self-consumption and vehicle-to-grid (V2G) for energy trading, adapting to real-time conditions through a hybrid long short-term memory (LSTM) neural network for accurate household load prediction, alongside a comprehensive nonlinear battery degradation model accounting for both cycle and calendar aging. Simulation results reveal significant economic advantages: compared to smart unidirectional charging, the proposed method yields an annual economic benefit of up to €3046.81, despite a modest 1.96% increase in battery degradation. Even under unfavorable market conditions, where V2G energy selling generates no revenue, V2H alone ensures yearly savings of €425.48. A systematic sensitivity analysis investigates how variations in battery capacity, household load, and price ratios affect economic outcomes, confirming the consistent benefits of bidirectional energy exchange. These findings highlight the potential of EVs as active energy nodes, enabling sustainable energy management and cost-effective battery usage in real-world conditions.
I INTRODUCTION
@@
In recent years, the widespread adoption of electric vehicles (EVs) has experienced a significant growth, bringing major changes to the transportation sector. EVs are providing a more sustainable solution to the challenges posed by climate change. However, this transition is also having a significant impact on global energy demand. With the exponential growth in the number of EVs on the road, the need for energy to recharge their batteries increases daily, placing significant challenges on electric grids. Projections indicate that global electricity demand for EV charging is expected to increase by 633% from 2023 to 2030 [1].
@@
The primary goal of traditional charging systems is to charge the battery once an EV is connected to the power socket [2]. However, a vehicle is parked approximately 96% of the time [3]. This highlights the potential to leverage EVs for active interaction with electric grids. With the emergence of bidirectional power transfer for EVs, it is now possible to reduce the costs for EV owners while providing a service to the grid. In this context, concepts such as vehicle-to-grid (V2G) and vehicle-to-home (V2H) are emerging as groundbreaking solutions, where EVs are no longer merely means of transportation but actual energy nodes. V2G allows EVs to supply energy back to the grid, providing services such as load balancing and frequency regulation [4], while allowing the EV owner to generate profit by selling stored energy. V2H allows EVs to supply power to a home, supporting home energy management and enabling the user to reduce energy costs and increase self-sufficiency [5]. In addition to being used to minimize electricity consumption costs, the V2H can also be used as a backup source for the load connected to the home in the event of a power failure from the grid [4].
The adoption of V2G and V2H technologies largely depends on user acceptance and preferences. A survey conducted in Sweden highlights that vehicle-home-grid integration presents both challenges and opportunities [6]. First, users tend to prefer V2H systems, perceiving them as relatively straightforward. This preference is primarily driven by the immediate financial benefits (i.e., energy cost savings for household consumption) and the added resilience provided during grid outages. While V2G is also seen as promising for its various services to the grid and the offered opportunity for energy arbitrage, this technology presents great challenges, particularly regarding battery longevity and the risk of insufficient charge capacity for daily vehicle use.
The user preferences highlighted in [6] underline the importance of addressing both technical and economic challenges to enhance the adoption of bidirectional charging systems. To this end, considerable research effort has been devoted to developing energy management algorithms for V2G and V2H.
@@
In the relevant literature, most approaches rely on offline optimization. For example, a multi-integer linear programming (MILP) framework was developed to optimize the operation of smart households using V2G and V2H, aiming to minimize user costs [7]. Without considering V2H, Dong et al. formulated the day-ahead V2G scheduling problem within the multi-agent reinforcement learning framework to optimize the peak shaving performance for the electric grid [8]. V2G and V2H applications for energy trading were also explored in [9], which reports that an average German household with a photovoltaic system, heat pump, and stationary battery can generate annual revenues of approximately €310. However, these three works ignored battery degradation in the optimization for simplicity, which tend to result in suboptimal solutions due to excessive battery costs. To mitigate the risk, Khezri et al. presented a MILP-based optimal scheduling model for EVs, in which V2G and a linearized battery degradation model were considered [10]. Following this trend, Lee et al. [11] presented a linear programming-based V2G optimization focusing on frequency regulation according to the rules of the Swedish market. A common feature of these works is that they assume full knowledge of input data in advance, such as household loads, and then solve the problem offline, making them less applicable to real-world with dynamic scenarios.
@@
To mitigate the potential drawbacks of offline optimization, recent research has increasingly focused on optimizing V2G technology online. Li et al. [12] proposed an online battery anti-aging V2G scheduling method and utilized fuzzy logic control to solve the optimization in two stages. Therein, only the cycle aging was considered during offline calibration to set the fuzzy control parameters, whereas battery degradation attributed to both cycle aging and calendar aging was not explicitly considered during the online phase. In [13], particle swarm optimization (PSO) was used to solve the V2G optimization with a sliding window; as for the battery degradation, it only considers the cycle aging with the rain-flow cycle counting method. Although battery degradation has been considered, the existing V2G optimization algorithms heavily rely overly simplified degradation models that fail to capture calendar and cycle aging comprehensively. Furthermore, these online algorithms have not systematically synergized V2G, V2H, battery dynamics, and the inherent uncertainties associated with household energy consumption. These limitations highlight the need for more practical and accurate solutions that can adapt to real-time conditions while mitigating the impact on battery lifespan.
@@
To address the identified research gaps, this paper proposes a nonlinear online optimization algorithm for the vehicle-home-grid integration, aiming to minimize user costs through energy trading based on hourly price variations. This algorithm is carried out by considering a single user, who owns both an EV and a house. To obtain more predictable results in real-world scenarios, the algorithm is implemented in an online manner, meaning that data is processed as it arrives rather than being known in advance. Specifically, we introduce a hybrid long short-term memory (LSTM) neural network to predict the domestic load, enhancing real-time energy allocation. Additionally, we use a detailed battery model to account for both calendar and cycle aging, particularly the nonlinear aging characteristics in response to various coupled stress factors. As a result, the proposed algorithm minimizes the overall user electricity and battery costs, contributing to more sustainable and economically viable energy management.
II METHODOLOGY
@@
The considered vehicle-home-grid integration, as illustrated in Fig. 1, has three actors: an EV, a house, and the power grid. The EV can supply energy to the grid and the house (V2G and V2H), while grid-to-vehicle (G2V) represents the energy used to recharge the EV. The house, on the other hand, can also receive energy in a standard way, directly from the power grid through grid-to-home (G2H).
Figure 1: Vehicle-home-grid interaction of the proposed algorithm. The arrows indicate the allowable energy flows for V2H, V2G, G2V and G2H.
As most EV users prefer to charge their EV at home rather than at public charging points [14], the present work assumes that the EV is only charged when parked at home.
@@
The electricity price is known in advance, thanks to the day-ahead pricing. However, the household load is only known up to the current time, while future values remain uncertain. To address this uncertainty, an online hybrid LSTM neural network is used to predict the future household load. Additionally, at each time step a correction algorithm refines the predicted values based on actual consumption, making it possible for the control algorithm to calculate the optimal energy flows within the given time window.
II-A Vehicle-Home-Grid Control
@@
$\min\sum_{t}EC_{t}+BC_{t}+s_{t},$ (1)
where $t$ represents time, $s_{t}$ is a slack variable for a soft constraint related to the EV’s SoC at the pickup time, and $EC_{t}$ and $BC_{t}$ denote the energy and battery costs, respectively.
@@
Based on economic principles discussed in [15], $NV$ is calculated through
@@
$SoC_{t}=SoC_{t-1}+\frac{E^{G2V}_{t}}{E_{b}}-\>$ $\frac{E^{V2G}_{t}}{E_{b}}$
$-\frac{E^{V2H}_{t}}{E_{b}}-\frac{D}{D_{r}},\>\>\>\>$ (9)
$SoC_{t}+s_{t}\geq\>$ $SoC^{goal}_{t},$ (10)
@@
In (5), all energy flows and $s_{t}$ must be non-negative. The input energy $E^{G2V}_{t}$ and the output energy from the EV ($E^{V2G}_{t}+E^{V2H}_{t}$) cannot exceed the maximum limit $E_{\max}$, imposed by the EV charger, resulting in the constraints formulated in (6) and (7). The constraint (8) ensures that the EV’s SoC ranges from 0 to 100%. The SoC evolution in (9) depends on the previous SoC, the energy exchanged with the battery (normalized by its capacity, $E_{b}$), and the distance traveled (normalized by the battery range, $D_{r}$). Notice that, while the EV is parked, $D=0$. (10) is a soft constraint ensuring the EV reaches the desired SoC of the user. Specifically, it requires the SoC to be at least $SoC^{goal}_{t}$, which is 80% at pickup times and zero otherwise. If the constraint is not met, $s_{t}$ compensates for the deviation, contributing to an increase in the objective function (1). The household load $HL_{t}$ is met by energy supplied either from the grid or the EV, according to (11). However, the total energy purchased from the grid cannot exceed the available supply $G_{t}$ (which is here assumed to be sufficiently large to always ensure the constraint (12) is never active).
Note that (1)–(12) apply only when the EV is parked. However, the model simulates a realistic scenario where the EV alternates between parking and driving throughout the simulation. During driving, no optimization is needed, as the EV is absent and all its energy flows are zero. In this phase, the SoC depends on the previous step and the distance traveled $D$, while the household load is fully supplied by the grid ($E_{t}^{G2H}$).
II-B Battery Model
An empirical battery model is employed to describe the degradation characteristics, including both calendar and cycle aging. This model has been experimentally validated against real-world lithium-iron-phosphate batteries [16] and has been widely applied in the literature [17]. While the details of the model can be found in [16], we present the key equations here for calculating the battery degradation in the vehicle-home-grid control.
@@ II-B Battery Model
where $U_{a}$ is the anode open-circuit potential (computed as a function of SoC) [16]. We assume a constant $T$ over time, equal to 15°C.
@@
Using the LiFePO4 battery model, the cell open-circuit voltage from [16] is scaled to a nominal voltage of 400 V, as shown in Fig. 2. Since the control algorithm operates in discrete time with a sampling time of one hour, the voltage transients are not considered. Instead, at each time step $t$, we approximate the voltage value by the average between $V(SoC_{t-1})$ and $V(SoC_{t})$.
Figure 2: Voltage profile with respect to SoC for the battery pack.
II-C Household Load Prediction and Management
In real-world applications, the household load is largely affected by weather and user behavior. Due to substantial uncertainties and stochasticity, the load values are typically unknown in advance. This poses challenges for the optimal energy flow to be computed when the EV is parked. To address this, we develop a real-time household load predictor based on historical data and patterns. Specifically, a hybrid long short-term memory neural network (LSTM) is applied. LSTM is very powerful in capturing dependencies in sequential data and is particularly suitable for time series energy consumption data that exhibit recurring daily and seasonal patterns. The hybrid LSTM approach is adopted from [18, 19] to incorporate additional features that influence energy consumption, providing the model with more contextual information and improved prediction accuracy.
@@ 1. $f1)$
    energy consumption from the previous 23 hours up to the current time step (from $t-23$ to $t$), which serves as the input sequence for the LSTM neural network
@@ 3. $f4)$
The architecture used, extensively studied in [19], is summarized in Fig. 3.
Figure 3: Hybrid LSTM neural network architecture, where nn labels the number of neurons.
@@
The hybrid LSTM neural network was implemented using Tensorflow and Keras. The model was trained with a batch size of 8 and for 75 epochs. The model employed the Adam optimizer. The other settings are shown in Fig. 3. As demonstrated in [19], this architecture shows accurate performances for electrical load prediction.
The proposed prediction model works during the time window in which the EV is parked. Based on the household load predictions, the optimization problem with the objective function in (1) is solved to determine the optimal energy profiles. However, as time progresses, real consumption data become available for the earliest predicted hour. This allows for online energy profile corrections in case the prediction deviates from the actual household load. The main steps of this adaptive optimization are given in Algorithm 1. $t_{arr}$ and $t_{pick}$ are the EV arrival time and the pickup time, respectively. $HL_{p}$ is the predicted household load.
Algorithm 1 Energy optimization during parking mode
1:Start the parking mode ($t=t_{arr}$)
2:while $t<t_{pick}$ do
3:     if $t=t_{arr}$ OR $HL(t)\neq HL_{p}(t)$ then
4:         PREDICT $HL$ using the hybrid LSTM struc-
5:          tured in Fig. 3 (from $t+1$ to $t_{pick}-1$)
6:         OPTIMIZE $E^{V2G}$, $E^{V2H}$, $E^{G2H}$, and $E_{t}^{G2V}$
7:         (from $t$ to $t_{pick}-1$)
8:     $t=t+1$
9:End the parking mode ($t=t_{pick}$)
Note that in Algorithm 1, the condition of $HL(t)\neq HL_{p}(t)$ is used to adjust the energy flow based on household load predictions. Specifically, if the predicted household load deviates from the actual consumption, both the prediction (Step 4) and optimization (Step 5) are updated accordingly. In an ideal scenario with a perfect load predictor, this adjustment would not be necessary, and the optimization would be performed only once within the algorithm for each parking session.
II-D Data Sources and Simulation Data Generation
For the household load, the dataset from [20] has been used in this work. It contains household load data for single apartments in the US, each spanning one year. Five datasets from the state of Washington have been selected because they exhibit similar consumption patterns (shown in Fig. 4, where the datasets are concatenated to form a five-year dataset).
@@ II-D Data Sources and Simulation Data Generation
Four of the five datasets were used for training, while one was used for testing, meaning 20% of the data was reserved for the model simulation. This last year of data, dedicated for the simulations, has an average hourly consumption of 0.9 kWh, resulting in a daily average consumption of 21.6 kWh.
@@
The electricity price for the year 2022 for Sweden (SE3 area) is taken from ENTSO-E [21]. However, this represents the initial price ($p^{ini}$), i.e., the price without taxes. The final price $p$ used in the simulations has been calculated using
@@
A 25% markup is applied to the initial price to approximately account for Swedish value-added tax (VAT), and a fixed cost of 0.006 €/kWh is added to represent the energy tax.
To match the resolution of the energy price and household load datasets, simulations are conducted over one year with an hourly time step. The simulation started with the battery with an age of 60 days and the EV’s SoC at 60%. The complete control algorithm was developed in Python, with CasADi [22] serving as the optimization framework and the IPOPT solver applied for solving the nonlinear optimization problems.
III RESULTS AND DISCUSSION

### v3

# Online Aging-Aware Energy Optimization for Vehicle-Home-Grid Integration

@@ Abstract
This paper investigates the economic impact of vehicle-home-grid integration through an online optimization algorithm that manages energy flows between an electric vehicle, a household, and the electrical grid. The algorithm exploits vehicle-to-home (V2H) for self-consumption and vehicle-to-grid (V2G) for energy trading, adapting in real-time via a hybrid long short-term memory (LSTM) network for household load prediction and a nonlinear battery degradation model including cycle and calendar aging. Simulations show annual economic benefits up to €3046.81 compared to smart unidirectional charging, despite a modest 1.96% increase in battery aging. Even under unfavorable market conditions, with no V2G revenue, V2H alone provides yearly savings of €425.48. Sensitivity analyses on battery capacity, household load, and price ratios confirm the consistent benefits of bidirectional energy exchange, highlighting the role of EVs as active energy nodes for sustainable management.
keywords:
Vehicle-to-grid, Vehicle-to-home, Battery aging, Household load prediction.
thanks: This work was supported by the Swedish Energy Agency within the Vehicle Strategic Research and Innovation Program (Grant No. P2022-00960) and the European Union’s Horizon Europe program under the Marie Skłodowska-Curie (Grant No. 101131278).
1 INTRODUCTION
@@
In recent years, the widespread adoption of electric vehicles (EVs) has experienced a significant growth, bringing major changes to the transportation sector. EVs are providing a more sustainable solution to the challenges posed by climate change. However, this transition is also having a significant impact on global energy demand. With the exponential growth in the number of EVs on the road, the need for energy to recharge their batteries increases daily, placing significant challenges on electric grids.
@@
The primary goal of traditional charging systems is to charge the battery once an EV is connected to the power socket. However, a vehicle is parked approximately 96% of the time (Kou et al. (2015)). This highlights the potential to leverage EVs for active interaction with electric grids. With the emergence of bidirectional power transfer, it is possible to reduce the costs for EV owners while supporting grid services. In this context, concepts such as vehicle-to-grid (V2G) and vehicle-to-home (V2H) are redefining EVs from mere transportation means into active energy nodes. V2G allows EVs to supply energy back to the grid, providing services such as load balancing and frequency regulation, while allowing the EV owner to generate profit. V2H allows EVs to supply power to a home, supporting home energy management, reducing costs and enhancing self-sufficiency (Islam et al. (2022)).
@@
Motivated by these opportunities, recent research has focused on optimizing energy exchanges between EVs, households, and the grid. In the relevant literature, most approaches rely on offline optimization. For example, a mixed-integer linear programming (MILP) framework was developed by Erdinc et al. (2015) to optimize the operation of smart households using V2G and V2H, aiming to minimize user costs. V2G and V2H applications for energy trading were also explored by Kern et al. (2022), which reports that an average German household with a photovoltaic system, heat pump, and stationary battery can generate annual revenues of approximately €310. However, these two works ignored battery degradation in the optimization for simplicity, which tend to result in suboptimal solutions due to excessive battery costs. To mitigate the risk, Khezri et al. (2024) presented a MILP-based optimal scheduling model for EVs, in which V2G and a linearized battery degradation model were considered. Following this trend, Lee et al. (2023) presented a linear programming-based V2G optimization focusing on frequency regulation according to the rules of the Swedish market. A common feature of these works is that they assume full knowledge of input data in advance, and then solve the problem offline, making them less applicable to real-world with dynamic scenarios.
@@
To mitigate the potential drawbacks of offline optimization, recent research has increasingly focused on optimizing V2G technology online. Li et al. (2022) proposed an online V2G scheduling method based on fuzzy logic control to mitigate battery aging through a two-stage optimization. Cycle aging was considered only during offline calibration of the fuzzy parameters, while calendar aging was not explicitly accounted for during the online operation. Zhang et al. (2024) used a particle swarm optimization to solve the V2G optimization with a sliding window; as for the battery degradation, it only considers the cycle aging with the rain-flow cycle counting method. Although battery degradation has been considered, the existing V2G optimization algorithms heavily rely overly simplified degradation models that fail to capture calendar and cycle aging comprehensively. Furthermore, these online algorithms have not systematically synergized V2G, V2H, battery dynamics, and the inherent uncertainties associated with household energy consumption.
@@
To address the identified research gaps, this paper proposes a nonlinear online optimization algorithm for vehicle-home-grid integration, aiming to minimize user costs through energy trading based on hourly price variations. The framework considers a single user owning both an EV and a household. To operate online, the algorithm employs a hybrid long short-term memory (LSTM) neural network to forecast the household load, while the energy scheduling is determined through an online optimization procedure inspired by shrinking-horizon model predictive control (MPC). A detailed battery model is incorporated to capture both calendar and cycle aging, accounting for their nonlinear dependence on coupled stress factors. As a result, the proposed method minimizes overall electricity expenses and battery degradation costs, contributing to sustainable and economically efficient energy management.
2 METHODOLOGY
@@
The considered vehicle-home-grid integration, as illustrated in Fig. 1, has three actors: an EV, a house, and the power grid. The EV can supply energy to the grid and the house (V2G and V2H), while grid-to-vehicle (G2V) represents the energy used to recharge the EV. The house, on the other hand, can also receive energy in a standard way, directly from the power grid through grid-to-home (G2H). Therefore, the present work assumes that the EV is only charged when parked at home.
Figure 1: Proposed vehicle-home-grid integration framework.
@@
The electricity price is known in advance, thanks to the day-ahead pricing. However, the household load is only known up to the current time, while future values remain uncertain. To address this uncertainty, a hybrid LSTM neural network is used to predict the future household load. Additionally, at each time step the controller updates the predicted values using the newly available data.
The control strategy is formulated as a shrinking-horizon MPC problem, activated whenever the EV is parked. The prediction horizon corresponds to the parking duration, known only at the beginning of each parking session.
2.1 Vehicle-Home-Grid Control
@@
$\min\sum_{t}EC_{t}+BC_{t},$ (1)
where $t$ represents time, and $EC_{t}$ and $BC_{t}$ denote the energy and battery costs, respectively.
@@
Based on economic principles discussed in Park (2013), $NV$ is calculated through
@@
$SoC_{t}=SoC_{t-1}+\frac{E^{G2V}_{t}}{E_{b}}-\>\frac{E^{V2G}_{t}}{E_{b}}$ $-\frac{E^{V2H}_{t}}{E_{b}}-\frac{d_{t}}{D_{r}},\>\>$ (9)
$SoC_{t}\geq\>$ $SoC^{goal}_{t},$ (10)
@@
In (5), all energy flows must be non-negative. The input energy $E^{G2V}_{t}$ and the output energy from the EV ($E^{V2G}_{t}+E^{V2H}_{t}$) cannot exceed the maximum limit $E_{\max}$, imposed by the EV charger, resulting in the constraints formulated in (6) and (7). The constraint (8) ensures that the EV’s SoC ranges from 0 to 100%. The SoC evolution in (9) depends on the previous SoC, the energy exchanged with the battery (normalized by its capacity, $E_{b}$), and the distance traveled (normalized by the battery range, $D_{r}$). Notice that, while the EV is parked, $d_{t}=0$. (10) requires the SoC to be at least $SoC^{goal}_{t}$, which is 80% at pickup times and zero otherwise. The household load $HL_{t}$ is met by energy supplied either from the grid or the EV, according to (11). However, the total energy purchased from the grid cannot exceed the available supply $G_{t}$ (which is here assumed to be sufficiently large to always ensure the constraint (12) is never active).
Note that (1)–(12) apply only when the EV is parked. However, the model simulates a realistic scenario where the EV alternates between parking and driving throughout the simulation. During driving, no optimization is needed, as the EV is absent and all its energy flows are zero. In this phase, the SoC depends on the previous step and the distance traveled $d_{t}$, while the household load is fully supplied by the grid ($E_{t}^{G2H}$).
2.2 Battery Model
An empirical battery model, experimentally validated on real-world lithium-iron-phosphate cells, is employed to describe both calendar and cycle aging. The complete model details and parameters are provided by Schimpe et al. (2018), while we present only the key equations relevant to the vehicle-home-grid control.
@@ 2.2 Battery Model
where $U_{a}$ is the anode open-circuit potential (computed as a function of SoC). We assume a constant $T$ over time, equal to 15°C.
@@
2.3 Household Load Prediction and Management
In real-world applications, the household load is largely affected by weather and user behavior. Due to substantial uncertainties and stochasticity, the load values are typically unknown in advance. This poses challenges for the optimal energy flow to be computed when the EV is parked. To address this, we develop a real-time household load predictor based on historical data and patterns. Specifically, a hybrid long short-term memory neural network (LSTM) is applied. LSTM is very powerful in capturing dependencies in sequential data and is particularly suitable for time series energy consumption data that exhibit recurring daily and seasonal patterns. The hybrid LSTM approach is adopted from (Huang et al. (2021); Lu et al. (2024)) to incorporate additional features that influence energy consumption, providing the model with more contextual information and improved prediction accuracy.
@@ 1. $f1)$
    energy consumption from the previous 23 hours up to the current time step (from t23t-23 to $t$), which serves as the input sequence for the LSTM neural network
@@ 3. $f4)$
The architecture used, extensively studied in Lu et al. (2024), is summarized in Fig. 2.
Figure 2: Hybrid LSTM neural network architecture, where $n$ labels the number of neurons.
@@
The hybrid LSTM neural network was implemented using Tensorflow and Keras. The model was trained with a batch size of 8 and for 75 epochs. The model employed the Adam optimizer. The other settings are shown in Fig. 2. As demonstrated in Lu et al. (2024), this architecture shows accurate performances for electrical load prediction.
The proposed prediction framework operates during the time window in which the EV is parked. During this interval, the household load is forecasted, and the predicted values are used as deterministic inputs to the optimization problem in (1), which determines the energy scheduling for the upcoming hours. As time progresses, the actual household consumption becomes available, replacing the corresponding predicted values. The optimization is then re-solved with this updated information, effectively implementing a shrinking-horizon MPC scheme.
2.4 Data Sources and Simulation Data Generation
For the household load, the dataset from National Renewable Energy Laboratory (2021) has been used in this work. It contains household load data for single apartments in the US, each spanning one year. Five datasets from the state of Washington have been selected because they exhibit similar consumption patterns (shown in Fig. 3, where the datasets are concatenated to form a five-year dataset).
@@ 2.4 Data Sources and Simulation Data Generation
Four of the five datasets were used for training, while one was used for testing, meaning 20% of the data was reserved for the model simulation. This last year of data, dedicated for the simulations, has an average hourly use of 0.9 kWh, resulting in a daily average consumption of 21.6 kWh.
@@
The electricity price for the year 2022 for Sweden (SE3 area) is taken from ENTSO-E (2025). However, this represents the initial price ($p^{ini}$), i.e., the price without taxes. The final price $p$ used in the simulations has been calculated using
@@
A 25% markup is applied to the initial price to account for Swedish value-added tax (VAT), and a fixed cost of 0.006 €/kWh is added to represent the energy tax.
To match the resolution of the energy price and household load datasets, simulations are conducted over one year with an hourly time step. The simulation started with the battery with an age of 60 days and the EV’s SoC at 60%. The complete control algorithm was developed in Python, with CasADi (Andersson et al. (2019)) serving as the optimization framework and the IPOPT solver applied for solving the nonlinear optimization problems.
3 RESULTS AND DISCUSSION